OpenStudy (anonymous):

If n-2 = 1/16, then n could be which of the following? a. -8 b. -4 c. -1/4 d. 1/4 e. 4 select all..

OpenStudy (anonymous):

Add 2 both the sides..

OpenStudy (campbell_st):

add 2 to both sides for the solution...

mathslover (mathslover):

\[\large{n-2=\frac{1}{16}}\] \[\large{n=\frac{1}{16}+2}\] \[\large{n=\frac{1+32}{2}}\] What r u getting ?

OpenStudy (anonymous):

the question is n^-2 =1/16 sorry

OpenStudy (jiteshmeghwal9):